# High Gain Fusion Target Design using Generative Artificial Intelligence

### v2

@@ I Return to Topological Basics
The topological fundamentals of fusion target and drive design were developed by Taylor [1]. It is based on the strong conservation, that is invariance even with interaction with external systems, of helicity. The helicity 3-form [2] is the well known Chern-Simmons 3-form [3, 4] that leads to the calculation of topological indices. The 2D equivalent of helicity is vorticity. The conservation of helicity causes an inverse cascade, self-organization, and the emergence of force-free states. These states are only metastable and do disrupt, given enough time. The relaxation of the system to the metastable force-free state, as the weak Noether invariant, energy, is transferred to the external system, is called Taylor Relaxation.
@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [23]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [24]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\text{i}\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [25]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [26, 27]; or the Mayer Cluster Expansion [28]. This enables the unification of the four forces as a question of Lie group symmetries, that is a Grand Unified Theory (GUT) of fields, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [29] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the infinitesimal Lie group generator, for Einstein’s General Theory of Relativity [30], in 1958 [31].
@@ V Control of Collective Systems
Collectives are fields or swarms of elementary particles (making elementary fields), charged particles (making plasmas), molecules (making fluids), celestial bodies (making cosmoses), economic entities (making economies), persons (making societies) and so on. This new theory gives a way of controlling (that is, optimizing and stabilizing) collective systems, whose state is given, not by variables, but as a function, $f(x)$. This also can be viewed as topological characterization and control.

### v3

@@ I Return to Topological Basics
The topological fundamentals of fusion target and drive design were developed by Taylor [1]. It is based on the strong conservation, that is invariance even with interaction with external systems, of helicity. The helicity 3-form [2] is the well known Chern-Simons 3-form [3, 4] that leads to the calculation of topological indices. The 2D equivalent of helicity is vorticity. The conservation of helicity causes an inverse cascade, self-organization, and the emergence of force-free states. These states are only metastable and do disrupt, given enough time. The relaxation of the system to the metastable force-free state, as the weak Noether invariant, energy, is transferred to the external system, is called Taylor Relaxation.
@@ IV A Mathematically Logical Process of Renormalization
Dirac was very dissatisfied in his later life with the state of field theory [23]. Like Einstein, he felt that field theory should have a reality, and be a question, not of probability, but of geometry (manifold curvature and geodesic motion) or topology or group symmetry. He also felt that “renormalization is not a logical mathematical process”. The HST is a logical mathematical process of renormalization, enabling Heisenberg’s canonical approach to field theory, via calculation of the S-matrix. A Principle Components Analysis (PCA) of the HST gives: the renormalization spectrums, that is the solutions to the Renormalization Group Equations (RGEs); Heisenberg’s Scattering Matrix (S-Matrix) [24]; the $m$-Body Greens functions, which is obvious given the form of $G(z)=\text{i}\ln(R_{0}(z))$ in the formula for the HST, which is the well-known Greens function of a complex variable [25]; the $m$-body scattering cross sections; the Wigner-Weyl transformation [26, 27]; the Mayer Cluster Expansion [28]; or the functional Taylor expansion of the action functional. This enables the unification of the four forces as a question of Lie group symmetries, that is a Grand Unified Theory (GUT) of fields, or geodesic motion on the dynamical manifold. This can be seen by Murray Gell-Mann’s Standard Model [29] (where U(1) is the Lie group symmetry of the ElectroMagnetic force, SU(2) is the Lie group symmetry of the weak force, and SU(3) is the Lie group symmetry of the strong force); and Dirac writing down the Hamiltonian, that is the infinitesimal Lie group generator, for Einstein’s General Theory of Relativity [30], in 1958 [31].
@@ V Control of Collective Systems
Collectives are fields or swarms of elementary particles (making elementary fields), charged particles (making plasmas), molecules (making fluids), celestial bodies (making cosmoses), economic entities (making economies), persons (making societies) and so on. This new theory gives a way of controlling (that is, optimizing and stabilizing) collective systems, whose state is given, not by variables, but as a function, $f(x)$. This also can be viewed as topological characterization and control. The two-stage canonical auto-encoder is finding the analytic function; whose Riemann surface (that is, manifold), has the topology of the collective system; and whose non-stationary singularity spectrums are the Riemann moduli or topological indices (that is, the integrals of the Chern-Simons 3-form) of the collective system.